Calculate the total reversible thermodynamic work required to produce a spray of water of droplet diameter 2000 nm from 40 liters of water at

25C. Take the surface tension of water as 72 mN m size?  How many water drops would be produced, theoretically if the droplets are at all equal size?


this is all the information given in the text book.
Answer #1

Volume of one droplet = 4/3 π r3

= 4/3 × π ×(1000×10-9m)3 = 4.1888 ×10-18m3

Number of droplets = total volume / volume of 1 droplet

= 40 ×10-3 m3/(4.1888×10-18m3)

Number of water drops would be produced= 9.55×1015 (answer)

Total work required = surface tension × area

= 72×10-3N/m × 9.55×1015×(4×π×(1000×10-9m)2)

= 8640 N.m

= 8640 J. (Answer)

= 8.64 kJ (answer)